A student has carried out a moving boundary experiment using K2SO4 as an electrolyte. The data obtained was tabulated below:


Cation velocity: 11.43 mms-1

Cross-sectional area of tube: 28.56 mm2

Transference number of cation: 0.593

Anion concentration: 0.025 mol dm-3


Calculate the current flow for the boundary moves as far as 45 mm.


1
Expert's answer
2021-01-11T03:46:18-0500

"nk=\frac{ShCk}{mK}"


"mK=\frac{ShCk}{nk}=\frac{28.56*45*2.5*10^{-8}}{0.593}=0,00054"